OpenStudy (anonymous):

factor x^3+9x^2+26x+24. I got (x+1)(x+3)(x+2) am I right?

OpenStudy (anonymous):

no some mistake again... temme how u did and i'll post want went wrong buddy :)

OpenStudy (anonymous):

okay so first I noticed there would be the ( ) () (), then I put x's in all of them, then + and then (x+1)(x+3)(x+2)

OpenStudy (anonymous):

is it (x+1)(x+3)(x+6)?

OpenStudy (anonymous):

see basically, 24=product of the roots, so 24=-x1x2x3 26=(x1x2+x1x3+x2x3) 9=-(x1+x2+x3) we hav to start by guessing only...so let us substitue 1,-1,2,-2 (but intelligent guess will tell its -2) therefore x+2 is a root . then simplify and get other roots... :) the other roots are -3,-4

OpenStudy (anonymous):

(x+2)(x+4)(x+3) to check if the answer is rite, multiply our root 2*3*4=24 and 2+3+4=9 2*3+3*4+2*4=26 so done :)